# An Interpretable Deep-Learning Framework for Predicting Hospital Readmissions From Electronic Health Records

### v1

@@ 6 Conclusions
In this work we addressed the problem of predicting unplanned hospital readmission from patients’ EHRs. We defined a novel framework, named ConvLSTM1d, that outperforms existing systems while providing at the same time more interpretable results. Our model is based on a ConvLSTM architecture to implement a temporal dimension that allows to consider past events and the disease history of patients. We validated ConvLSTM1d on two predictive tasks, namely unplanned hospital readmission within 30 and 180 days, using data extracted from the medical database MIMIC. Our solution achieves better performance than both traditional machine-learning models (based on logistic regression and random forest) and Deepr, currently the best deep-learning model for readmission prediction. Additionally, our results represents, as far as we know, the first use of a ConvLSTM network with word embeddings in the context of healthcare, not considering segmentation tasks on medical images. Moreover, we presented and validated a visual method to improve the interpretability of our predictions, based on an ad-hoc model-dependent procedure.
In the future, we will consider the application of this study to a bigger dataset, because at the moment MIMIC III is the only large public database available containing detailed information on patients’ EHRs. Moreover, time intervals and ICD-9-CM codes related to diagnoses and procedures represent a small portion of the information contained in EHRs. More information may be used to improve the predictive performance: for instance, for patients in intensive care unit, we may consider the history of prescribed medications and laboratory tests. Currently, our solution is based on few layers; a more elaborate and deeper network may also improve performance, but at the cost of reducing the interpretability of the model. Experiments with more complex architectures can also be included in future work.
@@ Acknowledgments
We are grateful to Letizia Tanca for her help in the definition and revision of this paper. This work was supported, in part, by Science Foundation Ireland grant 13/RC/2094_P2 and co-funded under the European Regional Development Fund through the Southern & Eastern Regional Operational Programme to Lero – the Science Foundation Ireland Research Centre for Software.

### v2

@@ 6 Conclusions
In this work, we addressed the problem of predicting unplanned hospital readmission from patients’ EHRs. We defined a novel neural network framework, named ConvLSTM1d, that outperforms existing traditional machine learning models while providing at the same time interpretable results. Our model is based on a ConvLSTM architecture to implement a temporal dimension that allows to consider past events and the full disease history of the patient. We tested ConvLSTM1d on two unplanned hospital readmission predictive tasks, within 30 and 180 days, using EHR data extracted from the large medical database MIMIC III. Our solution achieves higher accuracy compared to both traditional machine learning (logistic regression and random forest) and Deepr, a CNN -based deep learning model for predicting hospital readmissions. Additionally, our results represents, as far as we know, the first use of a ConvLSTM network with word embeddings for hospital readmission prediction, in conjunction with EHR data. Finally, we presented and validated a visual method to improve the interpretability of our predictions, based on custom model-dependent approach.
In the future, we will consider the application of this study to a larger dataset. At the moment, MIMIC III is the only large public database available containing detailed information on patients’ EHRs. Moreover, time intervals and ICD-9-CM codes related to diagnoses and procedures represent a small portion of the information contained in EHRs: more information may be used to improve the predictive performance, such as the history of prescribed medications and laboratory tests for patients in intensive care units. Currently, our solution is based on few layers. While this simplicity allows for a much more interpretable solution, we recognize that a more elaborate and deeper network may improve the performance of the model. However, this may come at the expense of reducing the overall interpretability of the system. Therefore, experiments and comparisons with more complex architectures (e.g., Transformers) will be included in future work. In particular, to achieve the double goal of improving performance while also keeping computational effort low, we plan to experiment with lightweight Transformers architectures such as Performer [17] or Linformer [20], and with a ConvLSTM network integrated with BERT-based language models for word vector encoding. We plan to consider both fine-tuned medical models such as ClinicalBERT [54] or BioBERT [55], and distillation approaches for reducing the size of the model, like DistilBERT [56] and TinyBERT [57]. Therefore, we plan to benchmark our current solution against attention-fueled solutions, both in terms of performance and interpretability, by evaluating the critical trade-off between complexity, accuracy and transparency in deep learning models.
@@ Acknowledgments
This work was supported, in part, by Science Foundation Ireland grant 13/RC/2094_P2 and co-funded under the European Regional Development Fund through the Southern & Eastern Regional Operational Programme to Lero – the Science Foundation Ireland Research Centre for Software.